# Take Fake as Real: Realistic-like Robust Black-box Adversarial Attack to Evade AIGC Detection

### v1

@@ III-C Implementation
This subsection provides the specific implementation of R2BA, with the pseudocode shown in Algorithm 1. The main notations are summarized in Table I.
@@
$D$()($\cdot$)( ⋅ ) Victim detector with output as forgery probability
@@
Particle swarm initialization. First, the algorithm initializes the global parameters $N$, $K$, $W_{\text{min}}$, $W_{\text{max}}$, $C_{\text{p}}$, $C_{\text{total}}$. Then, to avoid weak intensity that leads to poor adversariality, or excessive intensity that causes image distortion, set the range for the post-processing intensity. We formulate the $k$-th particle swarm as $P^{k}=(p_{1}^{k},p_{2}^{k},\dots,p_{N}^{k})$.
@@
1) The position $Pos(p_{i}^{k})$: That is the post-processing fusion intensity of the particle, which is formulated as Eq. 1.
@@
where $k=1$ and these 8 parameters symbolize the intensity of the four post-processing, randomly generated by Eq. 2. In Eq. 2, $\tau\in\{\alpha,\beta,\phi,\sigma,\lambda_{x},\lambda_{y},\theta,\gamma\}$, $\tau_{\text{min}}$ and $\tau_{\text{max}}$ is the lower bound and upper bound on the corresponding parameter.
@@
3) The fitness $f_{i}^{k}$: We calculate the fitness of each particle by Eq. 5. $X_{Pos(p_{i}^{k})}$ denotes the post-processed image of $X$, and $Pos(p_{i}^{k})$ is the post-processing intensity.
@@
where $D(\cdot)$ is the victim detector with a fake probability output. In other words, a smaller fitness value indicates a lower fake probability, implying better adversariality.
4) The individual best $f_{i,\text{best}}$ and $Pos_{\text{best}}(p_{i})$: The individual best, which refers to the best position $Pos_{\text{best}}(p_{i})$ corresponding to its all-time best fitness $f_{i,\text{best}}$. In initialization, we set the individual best by Eq. 6 and Eq. 7. and define the particle with the smallest fitness as the global best $f_{\text{best}}$ and $Pos_{\text{best}}(P)$.
@@
Optimization iteration process. Now we have completed the initialization. There comes the optimization iteration. Before each iteration begins, we need to determine particle’s current position $Pos(p_{i}^{k})$ and its next velocity $V(p_{i}^{k+1})$, which are the key factors that influence the particle’s next position $Pos(p_{i}^{k+1})$. Here the former is known. Therefore, the focus is on how to update the velocity of the particles. The update of the particle’s velocity involves two aspects: inertia $w$ and the two best $Pos_{\text{best}}(p_{i})$ and $Pos_{\text{best}}(P)$. The following describes how to update the particle’s velocity and position respectively.
@@
The inertia weight $w$ control the extent to which a particle depends on its previous velocity during updates. By gradually decreasing the inertia weight, R2BA can take different convergence modes during the optimization: Early in the optimization, larger inertia weight accelerates global exploration, enhances the ability to explore the entire example domain space, and reduces the risk of local optima. As optimization progresses and inertia decreases, particles can more precisely converge on the global best, reducing unnecessary position shifts and improving both convergence speed and accuracy, while also enhancing the ability to cross the decision boundary.
In R2BA, using linear decay allows the particles to gradually and smoothly towards the global best solution. The formula for inertia decay is given in Eq. 8.
@@
$V_{\text{p}}$ and $V_{\text{{total}}}$ represent the influence of the individual best position and the global best position from $k$-th iteration on the particle’s velocity in $k+1$-th iteration, respectively.
@@
Next, we update the particle’s position based on the updated velocity $v_{i}^{k+1}$, the previous position $Pos(p_{i}^{k})$ and a random modification term $\varepsilon_{i}$ by Eq. 12 and Eq. 13. The particle velocity incorporates both individual best and global best experiences, which are the key factors driving the particle towards the decision boundary.
@@
$\varepsilon_{i}=\begin{cases}random(),&\text{with prob }\epsilon,\quad r\sim U(0,1),\\ 0,&\text{with prob }1-\epsilon\end{cases}$ (13)
@@
Since the updated positions may exceed the defined parameter range, we need to set constraints on the particles’ positions. The basic constraints are given in Eq. 14. For certain parameters, such as the radius and coordinates, we need to ensure that they are integers or odd numbers. These constraints are described in Eq. 15 and Eq. 16.
@@ IV-A Experimental Setups
Basic attacks on white box and black box. We compare four white-box attacks, including FGSM[38], MIFGSM[39], PGD[40] and StatAT[9], and three black-box attacks, including SimBA[41], Square[32] and BruSLe[16]. StatAT[9] is the state-of-the-art (SOTA) white-box post-processing-based attack, and BruSLe[16] is the SOTA black-box query-based attack. For the parameters of baseline attacks, we set the max perturbation magnitude to $20/255$ with the pixel range $[0,1]$, the iterations number to $20$, the max number of queries is $1000$ and use random initialization to start attacks.
Implementation details. About image size, except for attacks on CNNSpot[24] and ResNetND[25] with $256\times 256$ image inputs, all other models use center-cropped $224\times 224$ image inputs to match the suitable input size of the pretrained models. For better analysis, we use bold font to indicate the best result and underline the second best result. We set the max number of queries to 1000 and others are set as follows:
@@ IV-B Anti-detection Performance Analysis
Attacks on GAN-based detectors. The experiments in Table V prove that the attack success rate of R2BA reaches the SOTA level, with a max improvement of 37.7% in ASR and an average ASR improvement of about 10%. In terms of image quality, it is at the best level in all the other three detectors except UnivFD[10] and the average results, indicating the excellent invisibility of R2BA. Since we limited the max perturbation magnitude for gradient-based attacks, their image quality on UnivFD is slightly better than ours. However, the attack effectiveness of these white-box methods against the UnivFD detector is nearly zero, confirming that gradient-based white-box attacks struggle with effective anti-detection performance.
Attacks on Diffusion-based detectors. The results in Table VI show that, on GenImage[3] and DRCT[4], R2BA has a maximum improvement of 41.2% in ASR and an average ASR improvement of approximately 18%. Additionally, the image quality metric BRISQUE[42] of R2BA is at an best level, while SSIM[35] is also either close to best or indeed best, demonstrating the excellent invisibility of R2BA in attacking diffusion-based AIGC detectors.
Comprehensive analysis. The results of on three popular AIGC datasets show that perturbations generated by traditional gradient-based or random pixel-modification attacks[16],[38],[39],[40],[41] are often detected as fake by well-trained AIGC detectors. As a post-processing-based white-box attack, due to based on statistical feature consistency, StatAT[9] performs well on facial images but poorly on multi-class natural images, as this type of methods cannot unify the feature dimensions on natural images with inherent multi-class nature and complexity. While Square[32] performs better in ASR but worse in image quality metrics compared to other algorithms, which suggests it may improve the anti-detection performance at the expense of image quality.
In contrast, as a post-processing-based adversarial attack, R2BA does not require consideration of feature-level consistency. Instead, it targets the vulnerabilities of the detectors to achieve efficient attacks, making it an adversarial attack method that is independent of the image classes. The results indicate that, by optimizing the post-processing fusion intensity, R2BA achieves impressive and comprehensive anti-detection performance. It successfully evades both GAN-based and diffusion-based AIGC detectors. The realistic-like adversarial examples generated by R2BA, matching the pattern of the real-world images, lead the detectors to recognize them as real natural images. In other words, when dealing with post-processing fusion attacks, the existing AIGC detectors have significant performance degradation, which suggests that if the AIGC detectors are put into practical application, their security cannot be guaranteed and remains to be further explored.
@@ IV-C Robustness Comparison Experiments
The robustness of adversarial examples to common image post-processing techniques is crucial for determining their real-world applicability. If an adversarial example is vulnerable to such techniques, its adversariality is erased, and the threat to AIGC detectors is eliminated.
@@
Therefore, we further conduct robustness comparison experiments for these adversarial attacks. Since our method is designed based on four post-processing methods, it is naturally robust to common platform image post-processing. Table VII shows robustness experiments against known post-processing, JPEG compression and Gaussian noise, while Table VIII compares robustness against unknown post-processing, Rotation and Resize (not included in the four post-processing we selected), offering a more comprehensive analysis. AVG in both tables refers to average performance.
@@
Against known image post-processing. Apparently, The experimental results in Table VII show that R2BA far outperforms other adversarial attacks at most settings in robustness. This result indicates that the adversarial examples generated by R2BA can better maintain their adversariality in practical applications and have stronger realistic adaptability. This results also suggest the effectiveness of our post-processing fusion optimization method. Due to the post-processing fusion optimization used in the attack, the image with the best post-processing fusion intensity generated by R2BA, when subjected to another similar post-processing, will deviate from the best post-processing intensity, which theoretically leads to a certain degree of performance degradation in terms of anti-detection. And so it is. The robustness experiments in Table VII further confirm this, providing additional evidence for the effectiveness of optimization in R2BA.
@@
Against unknown image post-processing. In Table VIII, it can be seen that our method remains remarkably robust when dealing with the unknown post-processing. This is because, essentially, our proposed method, R2BA, is not a simple additive noise perturbation and will not have its adversariality erased by simple image post-processing techniques, such as rotation or resize. In contrast, apart from causing severe image distortion[32], adversarial attack methods based on simple additive noise perturbations are easily faded or eliminated after image post-processing, leading to attack failure.
@@
Overall, after image processing, although the anti-detection performance is some degraded, the ASR of R2BA remains around $0.5$ or even higher. And when a detector’s detection success rate is only 50%, its detection is unreliable. On the other hand, the robustness experiments demonstrate that stronger post-processing intensity is not always better. This is consistent with our discovery from the pre-experiments in Fig. 2 of subsection III-A. The adversarial examples generated with excessively strong post-processing intensity are heavily distorted, which in turn causes the detector to classify the post-processed AIGC image as fake. This also proves that, to some extent, the robustness of AIGC detectors to widely-used image post-processing techniques is uncertain and fluctuates, indicating a potential security risk.
@@ IV-D Visual Comparison and Analysis
The invisibility of the adversarial example is crucial in determining whether this perturbation is likely to be erased. Better visualisation increases the likelihood of this adversarial example being shared in real life or on the internet. Therefore, we provide a visual comparison in Fig. 4. Ten images from different classes (e.g., dogs, landscapes, cars, letters, etc.) are randomly chosen for the image quality comparison, and one additional image is selected for detailed comparison and perturbation visualization. Combined with the excellent anti-detection performance of R2BA analysed in subsection IV-B, we can observe that R2BA also maintains quite good image quality, demonstrates strong image realism, and achieves multi-class anti-detection performance on natural images.
It is worth noting that, to ensure objectivity and fairness, some images with obvious post-processing traces (e.g., visible light spot) are specially selected for comparison. From both the detailed comparison and perturbation visualization, our image quality is clearly superior. Moreover, the post-processed AIGC images have practical significance. For example, our light spot can simulate real-world flashlight, while JPEG compression and Gaussian blur mimic real-world image quality degradation. These modifications enhance the realism of adversarial examples, whether viewed from a human or a detector’s perspective. Additionally, the visualisation also shows that existing AIGC detectors are not robust in detecting post-processed AIGC images, and slight image post-processing operations may cause the detector to output incorrect results.
@@ IV-E Black-box Query Count Comparison Experiments
In practice, the number of queries to the detector is limited. A smaller number of queries can save time and control the budget. Therefore, as a black-box query-based attack, the number of queries is an important factor to consider. To ensure fairness, we check whether the adversarial point is reached after every 100 queries (we use 100 particles, where 1 particle represents 1 query). In Table IX, we use FatFormer[11] as the detector, which is the only one that covers all three datasets and generative models in its research and maintains good detection results for all 3 datasets. Evidently, the number of queries for R2BA is significantly smaller than for other black-box attacks. Therefore, compared to other query-based adversarial attacks, R2BA is able to maintain excellent anti-detection performance with fewer queries.
@@ IV-F Effect of Particle Number
To further explore the influence of particle number on R2BA optimization, we vary the number of particles and the number of iterations to compare the variations in in ASR, BRISQUE[42], SSIM[35], and the number of queries under different particle numbers. To ensure fairness of the experiments, the maximum number of queries is always kept at 1000.
The experiments are shown in Fig 5. We observe that when the number of particles increases, the ASR also increases. When the number of particles reaches 100, the ASR of all detectors reaches a quite high level. However, the effect of the particle number on BRISQUE and SSIM is not significant. This indicates that when the intensity of post-processing is controlled within a certain range, the image quality of the adversarial examples generated by R2BA does not fluctuate significantly with changes in other parameters of the algorithm.
Interestingly, in the experiment of UnivFD(DRCT), although we typically believe that an increase in the number of particles leads to a rise in the number of queries, the experimental results show that the number of queries is actually higher with a lower number of particles. Building on the previous experimental results regarding attack success rate, we know that UnivFD(DRCT) is a detector that is harder to attack. This suggests that when attacking robust detectors, our optimization algorithm struggles to find a suitable optimization direction with a lower number of particles, and thus more queries are needed for exploration. On the other hand, when the number of particles is higher, the query cost for a single iteration increases, and the increase in the query base leads to a multiplicative trend in the number of queries.
Therefore, choosing the right number of particles is crucial for optimizing the attack success rate of the algorithm. From the experimental results, it can be seen that choosing 100 particles and carrying out 10 iterations of the attack is a more ideal choice, as it achieves a better balance between the attack success rate and the number of queries.

### v2

@@ III-C Implementation
This subsection provides the specific implementation of R2BA, with the pseudocode presented in Algorithm 1. The main notations are summarized in Table I.
@@
$D$()($\cdot$)( ⋅ ) Victim detector with output as fake probability
@@
Particle swarm initialization. First, the algorithm initializes the global parameters $N$, $K$, $W_{\text{min}}$, $W_{\text{max}}$, $C_{\text{p}}$, $C_{\text{total}}$. Then, to avoid weak intensity, which leads to poor adversariality, or excessive intensity, which causes image distortion, the range for the post-processing intensity is set. We formulate the $k$-th particle swarm as $P^{k}=(p_{1}^{k},p_{2}^{k},\dots,p_{N}^{k})$.
@@
1) The position $Pos(p_{i}^{k})$: This is the post-processing fusion intensity of the particle, which is formulated as Eq. 1.
@@
where $k=1$ and these 8 parameters represent the intensity of the four post-processing, randomly generated by Eq. 2. In Eq. 2, $\tau\in\{\alpha,\beta,\phi,\sigma,\lambda_{x},\lambda_{y},\theta,\gamma\}$, $\tau_{\text{min}}$ and $\tau_{\text{max}}$ are the lower and upper bounds of the corresponding parameters.
@@
3) The fitness $f_{i}^{k}$: We calculate the fitness of each particle by Eq. 5. $X_{Pos(p_{i}^{k})}$ denotes the post-processed image of $X$, and $Pos(p_{i}^{k})$ is the post-processing fusion intensity.
@@
In other words, a smaller fitness value indicates a lower fake probability, implying better adversariality.
4) The individual best $f_{i,\text{best}}$ and $Pos_{\text{best}}(p_{i})$: The individual best refers to the best position $Pos_{\text{best}}(p_{i})$ corresponding to its all-time best fitness $f_{i,\text{best}}$. In initialization, we set the individual best by Eq. 6 and Eq. 7 and define the global best $f_{\text{best}}=\min\left\{f_{i,\text{best}}\right\}$ and $Pos_{\text{best}}(P)$ as the particle position corresponding to $f_{\text{best}}$.
@@
Optimization iteration process. Now we have completed the initialization. Here comes the optimization iteration. Before each iteration begins, we need to determine particle’s current position $Pos(p_{i}^{k})$ and its next velocity $V(p_{i}^{k+1})$, which are the key factors that influence the particle’s next position $Pos(p_{i}^{k+1})$. Here $Pos(p_{i}^{k})$ is known. Therefore, the focus is on how to update the velocity of the particles. The update of the particle’s velocity involves two aspects: inertia $w$ and the two best $Pos_{\text{best}}(p_{i})$ and $Pos_{\text{best}}(P)$. The following describes how to update the particle’s velocity and position respectively.
@@
The inertia weight $w$ controls how much a particle depends on its previous velocity during updates. By gradually decreasing the inertia weight, R2BA can take different convergence modes during the optimization: Early in the optimization, larger inertia weight accelerates global exploration, enhances the ability to explore the entire example domain space, and reduces the risk of local optima. As optimization progresses and inertia decreases, particles can more precisely converge on the global best, reducing unnecessary position shifts and improving both convergence speed and accuracy. In R2BA, using linear decay allows the particles to gradually and smoothly move towards the global best solution. The formula for inertia decay is given in Eq. 8.
@@
where $r_{\text{p}}$ and $r_{\text{total}}$ are random decimals in $[0,1)$, and $V_{\text{p}}$ and $V_{\text{{total}}}$ represent the influence of the individual best position and the global best position from $k$-th iteration on the particle’s velocity in $k+1$-th iteration, respectively.
@@
Next, we update the particle’s position based on the updated velocity $v_{i}^{k+1}$, the current position $Pos(p_{i}^{k})$ and a random modification term $\varepsilon_{i}$ by Eq. 12 and Eq. 13. The particle velocity incorporates both individual best and global best experiences, which are the key factors driving the particle towards the decision boundary.
@@
$\varepsilon_{i}=\begin{cases}random(),&\text{with prob }\epsilon,\\ 0,&\text{with prob }1-\epsilon\end{cases}$ (13)
@@
Since the updated position $Pos(p_{i}^{k+1})$ may exceed the defined parameter range, we need to set constraints on the particles’ positions. The basic constraints are given in Eq. 14. For certain parameters, such as the radius and coordinates, we need to ensure that they are integers or odd numbers. These constraints are described in Eq. 15 and Eq. 16.
@@ IV-A Experimental Setups
Basic attacks on white box and black box. We compare four white-box attacks, including FGSM[39], MIFGSM[40], PGD[41] and StatAT[9], and three black-box attacks, including SimBA[42], Square[32] and BruSLe[16]. StatAT[9] is the SOTA white-box post-processing-based attack, and BruSLe[16] is the SOTA black-box query-based attack. For the parameters of baseline attacks, we set the max perturbation magnitude to $20/255$ with the pixel range $[0,1]$, the iterations number to $20$, the max number of queries is $1000$ and use random initialization to start attacks.
Implementation details. About image size, except for attacks on CNNSpot[24] and ResNetND[25] with a $256\times 256$ image size, all other models use center-cropped $224\times 224$ images to match the suitable input size of the pretrained models. For better analyses, we use bold font to indicate the best result and underline the second best result. We set the max number of queries to 1000 and others are set as follows:
@@ IV-B Anti-detection Performance Analysis
Attacks on GAN-based detectors. The experiments in Table V prove that the attack success rate of R2BA reaches the SOTA level, with a max improvement of 38% in ASR and an average ASR improvement of about 10%. In terms of image quality, it is at the highest level among the average results and all GAN-based detectors except for the UnivFD[10] detector, indicating the excellent invisibility of our proposed method. Since we limited the max perturbation magnitude for gradient-based attacks, their image quality on UnivFD is slightly better than ours. However, the attack effectiveness of these white-box methods against the UnivFD detector is nearly zero, confirming that gradient-based white-box attacks struggle to achieve effective anti-detection performance.
Attacks on Diffusion-based detectors. The results in Table VI show that, on the GenImage[3] and DRCT[4] datasets, R2BA has a maximum improvement of 41% in ASR and an average ASR improvement of approximately 18%. Additionally, in terms of the average image quality, R2BA achieves an improvement of 33%–69% in BRISQUE[43] and 11%–77% in SSIM[36], indicating that R2BA demonstrates excellent invisibility when attacking diffusion-based AIGC detectors.
Comprehensive analysis. The results of on three popular AIGC datasets show that perturbations generated by traditional gradient-based or random pixel-modification attacks[16],[39],[40],[41],[42] are often detected as fake by well-trained AIGC detectors. As a post-processing-based white-box attack, due to based on statistical feature consistency, StatAT[9] performs well on facial images but poorly on multi-class natural images, as this type of methods cannot unify the feature dimensions on natural images with inherent multi-class nature and complexity. While Square[32] performs better in ASR, it performs worse in image quality metrics compared to other algorithms, suggesting that it may improve anti-detection performance at the expense of image quality.
In contrast, as a post-processing fusion adversarial attack, R2BA does not require consideration of feature-level consistency. Instead, it targets the vulnerabilities of the detectors to achieve efficient attacks, making it an image-class-independent adversarial attack. By optimizing the post-processing fusion intensity, R2BA achieves an average anti-detection performance improvement of 15%–72%, along with a significant improvement of 26%–68% in the BRISQUE[43] across the three datasets. It successfully evades both GAN-based and diffusion-based AIGC detectors. The realistic-like adversarial examples generated by R2BA, which match the pattern of the real-world images, cause the detectors to classify them as real natural images. In other words, when dealing with post-processing fusion attacks, the existing AIGC detectors suffer significant performance degradation, which suggests that if the AIGC detectors are put into practical application, their security cannot be guaranteed and requires further investigation.
@@ IV-C Robustness Comparison Experiments
The robustness of adversarial examples to common image post-processing techniques is crucial for their real-world applicability. If an adversarial example is vulnerable to such techniques, its adversariality is erased, and the threat to AIGC detectors is eliminated.
@@
Therefore, we further conduct robustness comparison experiments for these adversarial attacks. Since our method is based on four post-processing methods, it is naturally robust to common image post-processing. Table VII shows robustness experiments against known post-processing techniques, including JPEG compression and Gaussian noise, while Table VIII compares robustness against unknown post-processing techniques, including Rotation and Resize (not part of the four post-processing we select), offering a more comprehensive analysis. AVG in both tables refers to average performance.
@@
Against known image post-processing. Apparently, the experimental results in Table VII show that R2BA far outperforms other adversarial attacks in most settings for robustness. The results indicate that the adversarial examples generated by R2BA can better maintain their adversariality in practical applications and have stronger adaptability to real-world scenarios. Additionally, the results suggest the effectiveness of our post-processing fusion optimization method. Due to the post-processing fusion optimization used in the attack, the image with the best fusion intensity generated by R2BA, when subjected to another similar post-processing, will deviate from the best post-processing intensity, which theoretically leads to some degree of performance degradation in terms of anti-detection. This is indeed the case. The robustness experiments in Table VII further confirm this, providing additional evidence for the effectiveness of the optimization in R2BA.
@@
Against unknown image post-processing. In Table VIII, it can be seen that our method remains remarkably robust when dealing with unknown post-processing. This is because, essentially, our proposed method, R2BA, is not a simple additive noise perturbation and will not have its adversariality erased by simple image post-processing techniques, such as rotation or resizing. In contrast, apart from causing severe image distortion[32], adversarial attack methods based on simple additive noise perturbations are easily faded or eliminated after image post-processing, leading to attack failure.
@@
Overall, after image processing, although the anti-detection performance is somewhat degraded, the ASR metric of R2BA still shows a 21%–47% improvement compared to existing adversarial attack methods. Moreover, the ASR of R2BA in robust scenarios remains around 0.5 or even higher, suggesting that the reliability of the detector remains questionable. On the other hand, the robustness experiments demonstrate that stronger post-processing fusion intensity is not always better. This is consistent with our discovery from the pre-experiments in Fig. 2 of subsection III-A. Adversarial examples generated with excessively strong post-processing intensity are heavily distorted, which in turn causes the detector to classify the post-processed AIGC image as fake. This also proves that, to some extent, the robustness of AIGC detectors to widely-used image post-processing techniques is uncertain and fluctuates, indicating a potential security risk.
@@ IV-D Visual Comparison and Analysis
The invisibility of the adversarial example is crucial in determining whether this perturbation is likely to be detected. Better visualization increases the likelihood of this adversarial example being shared in real life or on the internet. Therefore, we provide a visual comparison in Fig. 4. Ten images from different classes (e.g., dogs, landscapes, cars, letters, etc.) are randomly chosen for the image quality comparison, and one additional image is selected for detailed comparison and perturbation visualization. Combined with the excellent anti-detection performance of R2BA analyzed in subsection IV-B, we can observe that R2BA also maintains high image quality, demonstrates strong image realism, and achieves multi-class anti-detection performance on natural images.
It is worth noting that, to ensure objectivity and fairness, some images with obvious post-processing traces (e.g., visible light spot) are specially selected for comparison. From both the detailed comparison and perturbation visualization, our image quality is clearly superior. Moreover, the post-processed AIGC images have practical significance. For example, our light spot can simulate a real-world flashlight, while JPEG compression and Gaussian blur mimic real-world image quality degradation. These modifications enhance the realism of adversarial examples, whether viewed from a human or detector’s perspective. Additionally, the visualization also shows that existing AIGC detectors are not robust in detecting post-processed AIGC images, and slight image post-processing operations may cause the detector to output incorrect results.
@@ IV-E Black-box Query Count Comparison Experiments
In practice, the number of queries to the detector is limited. A smaller number of queries can save time and control the budget. Therefore, as a black-box query-based attack, the number of queries is an important factor to consider. To ensure fairness, we check whether the adversarial point is reached after every 100 queries (we use 100 particles, with each particle representing 1 query). In Table IX, we use FatFormer[11] as the detector, which is the only one that covers all three datasets and generative models in its research and maintains good detection results for all three datasets. Evidently, the number of queries for R2BA is significantly smaller than that for other black-box attacks. Therefore, compared to other query-based adversarial attacks, R2BA is able to maintain excellent anti-detection performance with fewer queries.
@@ IV-F Effect of Particle Number
To further explore the influence of the particle number on R2BA optimization, we vary the number of particles and the number of iterations to compare the variations in ASR, BRISQUE[43], SSIM[36], and the number of queries under different particle numbers. To ensure the fairness of the experiments, the maximum number of queries is kept at 1000.
The experiments are shown in Fig. 5. We observe that when the number of particles increases, the ASR also increases. When the number of particles reaches 100, the ASR of all detectors reaches a relatively high level. However, the effect of the particle number on BRISQUE and SSIM is not significant. This indicates that when the intensity of post-processing is controlled within a certain range, the image quality of the adversarial examples generated by R2BA does not fluctuate significantly with changes in other parameters of R2BA.
Interestingly, in the UnivFD(DRCT) experiment, although we typically believe that an increase in the number of particles leads to a rise in the number of queries, the experimental results show that the number of queries is actually higher with a lower number of particles. Building on the previous experimental results regarding the attack success rate, we know that UnivFD(DRCT) is a detector that is harder to attack. This suggests that when attacking robust detectors, our optimization algorithm struggles to find a suitable direction for optimization with a lower number of particles, and thus more queries are needed for exploration. On the other hand, when the number of particles is higher, the query cost for a single iteration increases, and the increase in the query base leads to a multiplicative trend in the number of queries.
Therefore, choosing the right number of particles is crucial for optimizing the attack success rate of the algorithm. From the experimental results, it can be seen that choosing 100 particles and carrying out 10 iterations of the attack is the most ideal choice, as it strikes a better balance between the attack success rate and the number of queries.